# Fixing confirmation bias in feature attribution methods via semantic match

### 21 Feb 2024 v2

@@ 3.2 Testing hypothesis
Now that we are able to specify a hypothesis, the natural question to ask is: is this hypothesis valid? Zhou et al. [2022] define validity to be the expected proportion of tuples satisfying $A$ whose behavior complies to $B$ [^2] Additionally, Zhou et al. [2022] define coverage as $P(\bm{u}_{i}\models A_{\theta})$, namely the proportion of tuples triggering the hypothesis and sharpness as $P(\bm{u}_{i}\models A_{\theta}|M(\bm{u}_{i})\models B_{\theta})$.:
@@ 3.3 Experimental setup
SQuAD dataset. We also include an experiment on a common NLP task, question answering, on the SQuAD dataset. Ko et al. [2021] showed that it is possible to build biased models which focus only on specific parts of the context, by training them on data subsets that only contain answers in the $n$th sentence of the context. With this experiment, we want to employ semantic match to discriminate between a biased and an unbiased model. We investigate bias in the first sentence, as the subdataset containing only answers in the first sentence is the largest. Using a similar pipeline as Ko et al. [2021], we fine-tune two BERT-base uncased models [Devlin et al., 2019]; one on a biased dataset where the answer is always in the first sentence, and one on a dataset with approximately the same number of data points but with answers spread evenly among sentences. Details on implementation, as well as evidence on the fact that the first model is indeed biased, can be found in Appendix C. In this task, the BERT model outputs two independent probability distributions over the tokens in the context: one for the starting position and one for the ending position of the answer. Thus, we have SHAP values for both positions. We map each token in a sample to the corresponding sentence using spaCy sentencizer spa . Analogous to the image experiment, one can think of the (pre-image of the) mapping as drawing a bounding box around each sentence. We process the SHAP values to obtain the percentage of contribution of each sentence for the prediction (see Appendix C).
@@ Appendix C Further details and results on the SQuAD experiments
After training, we compute the SHAP values for the biased and unbiased model on the validation set. For each token of the answer we obtain a contribution vector with $n$ entries, where $n$ is the total number of sentences in the context; [^3] We exclude the contributions from the question, as we expect them to be relevant for the prediction in both models. Re-running the experiment including the question we obtain similar results. Each entry is the normalized sum of the contributions of each token in that sentence to the final SHAP value of the answer token. We do this both for starting and ending values. For each sentence (i.e., each entry in the contribution vector) we compute the maximum and average between the starting and ending contribution (in results we use the average contribution). The model can either return as prediction a single token (as in Figure 11) or multiple tokens. In the latter case we perform the above passages for each token of the prediction and then average across the tokens.
While computing the values, we find a bug in the SHAP library: the masking function used to evaluate the importance of features sometimes malfunctions when proper names appear in the question due to inconsistent tokenization, producing a longer output which in turn throws a length mismatch error. As the number of data points producing this error is around 10% of the total validation dataset, we exclude them from the experiment, listing their ids in the code repository for reproducibility.444 github.com/anonymized-for-submission To generate the mapping between tokens and sentences we use the SpaCy sentencizer of the en_core_web_lg model by [Honnibal and Montani, 2017].
@@
For hypothesis $\theta_{3}$ we expect something different. This is a ’sanity check’ hypothesis: we expect both models to put large contribution on the first sentence when the answer is there and they predict correctly. Hence if both pass this check we should see similar plots on both sides. The median distance plot of the third hypothesis (Figure 15) is similar to the previous two, essentially confirming that for the biased model the contribution of the first sentence is clustered around 50%. Contrary to the other hypotheses, now the unbiased model also reproduces the same behavior. The AUC plot for the biased model is also similar to the one obtained for $\theta_{2}$, and again the unbiased model matches it (the numbers are not identical, but the trend is comparable). We interpret these results as evidence that the contribution of the first sentence for the unbiased model is high only when the answer is in that sentence. For the biased model, conversely, the first sentence’s contribution is high regardless of the position of the true answer.

### 26 Feb 2024 v3

@@ 3.2 Testing hypothesis
Now that we are able to specify a hypothesis, the natural question to ask is: is this hypothesis valid? Zhou et al. [2022] define validity to be the expected proportion of tuples satisfying $A$ whose behavior complies to $B$: [^2] Additionally, Zhou et al. [2022] define coverage as $P(\bm{u}_{i}\models A_{\theta})$, namely the proportion of tuples triggering the hypothesis and sharpness as $P(\bm{u}_{i}\models A_{\theta}|M(\bm{u}_{i})\models B_{\theta})$.
@@ 3.3 Experimental setup
SQuAD dataset. We also include an experiment on a common NLP task, question answering, on the SQuAD dataset. Ko et al. [2021] showed that it is possible to build biased models which focus only on specific parts of the context, by training them on data subsets that only contain answers in the $n$th sentence of the context. With this experiment, we want to employ semantic match to discriminate between a biased and an unbiased model. We investigate bias in the first sentence, as the subdataset containing only answers in the first sentence is the largest. Using a similar pipeline as Ko et al. [2021], we fine-tune two BERT-base uncased models [Devlin et al., 2019]; one on a biased dataset where the answer is always in the first sentence, and one on a dataset with approximately the same number of data points but with answers spread evenly among sentences. Details on implementation, as well as evidence on the fact that the first model is indeed biased, can be found in Appendix C. In this task, the BERT model outputs two independent probability distributions over the tokens in the context: one for the starting position and one for the ending position of the answer. Thus, we have SHAP values for both positions. We map each token in a sample to the corresponding sentence using spaCy sentencizer [Honnibal et al., 2020]. Analogous to the image experiment, one can think of the (pre-image of the) mapping as drawing a bounding box around each sentence. We process the SHAP values to obtain the percentage of contribution of each sentence for the prediction (see Appendix C).
@@ Appendix C Further details and results on the SQuAD experiments
After training, we compute the SHAP values for the biased and unbiased model on the validation set. For each token of the answer we obtain a contribution vector with $n$ entries, where $n$ is the total number of sentences in the context; [^3] We exclude the contributions from the question, as we expect them to be relevant for the prediction in both models. Re-running the experiment including the question we obtain similar results. Each entry is the normalized sum of the contributions of each token in that sentence to the final SHAP value of the answer token. We do this both for starting and ending values. For each sentence (i.e., each entry in the contribution vector) we compute the maximum and average between the starting and ending contribution (in our results we use the average contribution). The model can either return as prediction a single token (as in Figure 11) or multiple tokens. In the latter case we perform the above passages for each token of the prediction and then average across the tokens.
While computing the values, we find a bug in the SHAP library: the masking function used to evaluate the importance of features sometimes malfunctions when proper names appear in the question due to inconsistent tokenization, producing a longer output which in turn throws a length mismatch error. As the number of data points producing this error is around 10% of the total validation dataset, we exclude them from the experiment, listing their ids in the code repository for reproducibility.444 github.com/anonymized-for-submission To generate the mapping between tokens and sentences we use the SpaCy sentencizer of the en_core_web_lg model [Honnibal et al., 2020].
@@
For hypothesis $\theta_{3}$ we expect something different. This is a ‘sanity check’ hypothesis: we expect both models to put large contribution on the first sentence when the answer is there and they predict correctly. Hence if both pass this check we should see similar plots on both sides. The median distance plot of the third hypothesis (Figure 15) is similar to the previous two, essentially confirming that for the biased model the contribution of the first sentence is clustered around 50%. Contrary to the other hypotheses, now the unbiased model also reproduces the same behavior. The AUC plot for the biased model is also similar to the one obtained for $\theta_{2}$, and again the unbiased model matches it (the numbers are not identical, but the trend is comparable). We interpret these results as evidence that the contribution of the first sentence for the unbiased model is high only when the answer is in that sentence. For the biased model, conversely, the first sentence’s contribution is high regardless of the position of the true answer.